20 g of ammonium nitrate (NH4NO3) is dissolved in 125 g of water in a coffee-cup calorimeter, the temperature of which falls from 296.5 K to 286.4 K. Find the  value of q for the calorimeter (Treat heat capacity of water as the heat capacity of the calorimeter and its contents). (Ans= -5.285 kJ)
 
Please explain why in the answer mass is taken to be only 125 g but not 125+20 g

Grade:

1 Answers

Abhinav priye
5 years ago
Mass of water =125 g(given)
Change in temp =286.4-296.5=-10.1
Specific heat of water in liquid form is 4.18
Heat capacity=specific heat capacity*given mass
=4.18*125=522.5
Value of q=522.5*-10.1=5277.8=5.28kj
Value of q=522.5*-10=